OpenStudy (anonymous):

Determine if the graph is symmetric about the x-axis, the y-axis, or the origin: r = 2 cos 5θ

OpenStudy (anonymous):

could someone answer the fluttering question

OpenStudy (kendricklamar2014):

@Wendy.Ivette11714

OpenStudy (wendy.ivette11714):

Convert polar coordinates to euclidean coordinates and prove that x= \y\ (absolute value) then it is symmetrical about the x axis the conversion formulas are \[x=rcosθ\] \[y=rsinθ\]

OpenStudy (wendy.ivette11714):

I think it would be best for the origin symmetry to prove wheter or not the radius is the same when you take opposite sides, this case would be tha radius in 0 and 180, in 90 and 270, in 45 and 225, and in 135 and 315. If they are not the same value then it's because it's not symmetrical. I havent worked with polar coordinates that much but I believe that you can use 90 and -90 (-90 being 270), but as I said I'm not sure.

OpenStudy (irishboy123):

\(r = 2 \cos 5θ\) try subbing \(\theta = - \theta \) to see if it is symmetrical about the Cartesian x axis  then try subbing \((\pi - \theta)\) for \(\theta\) for symmetry about the Cartesian y axis